Lamonda Corp. uses a job order cost system. On April 1, the accounts had balances as shown in the T-accounts below: The follo5. Prepare Lamondas April income statement. Include any adjustment to Cost of Goods Sold needed to dispose of over-or underaEnd. Bal. End. Bal. Cost of Goods Sold Sales Revenue Beg. Bal. Beg. Bal. End. Bal. End. Bal. Nonmanufacturing Expenses Beg. BRequired: 1. & 2. Post the April transactions to the T-accounts and compute the balance in the accounts at the end of April.Required: 1. & 2. Post the April transactions to the T-accounts and compute the balance in the accounts at the end of April.Complete this question by entering your answers in the tabs below. Req 1 and 2 Req За Req 3b Req 4 Req 5 Prepare Lamondas coComplete this question by entering your answers in the tabs below. Req 1 and 2 Req 3a Reg 3b Reg 4 Reg 5 Prepare Lamondas Ap
Answer #1

Part 1 & 2

Raw materials inventory

Bal.

29700

b.

112300

a.

232970

Bal.

150370

Work in process inventory

Bal.

19900

f.

262750

b.

96300

c.

181525

e.

250504.50

Bal.

285479.50

112300-16000= 96300

224900-43375 =181525

181525*138%= 250504.50

Finished goods inventory

Bal.

123700

g.

324270

f.

262750

Bal.

62180

Manufacturing overhead

b.

16000

e.

250504.50

c.

43375

d.

23700

d.

10500

d.

19200

137729.50

Overapplied

Cost of goods sold

g.

324270

Bal.

324270

Sales revenue

h.

518000

Bal.

518000

Non-manufacturing expenses

d.

35600

d.

50350

Bal.

85950

Part 3 A

Overapplied overhead = applied overhead – actual overhead = 250504.50-(16000+43375+23700+10500+19200) = $137729.50

Part 3 B

Decrease

Due to overapplied overhead there is decrease in cost of goods sold

Part 4

Lamonda Corp.

Cost of goods manufactured report

For the month of April

Beginning raw materials inventory

29700

Plus: raw materials purchase

232970

Less: indirect materials

16000

Less: ending raw materials inventory

150370

Direct materials used

96300

Direct labor

181525

Manufacturing overhead applied

250504.50

Total current manufacturing costs

528329.50

Plus: beginning work in process inventory

19900

Less: ending work in process inventory

285480

Cost of goods manufactured

262749.50

Part 5

Lamonda Corp.

Income statement

For the month of April

Sales revenue

518000

Cost of goods sold

Beginning finished goods inventory

123700

Plus: cost of goods manufactured

262749.50

Less: ending finished goods inventory

62180

Unadjusted cost of goods sold

324269.50

Less: overapplied manufacturing overhead

137729.50

186540

Gross profit

331460

Selling and administrative expenses

85950

Net income from operations

$245510